OpenStudy (anonymous):

derived... f(x)=sin \left( x+1/2x-3 \right)

OpenStudy (anonymous):

\[f(x)=\sin \left( x+1/2x-3 \right) \]

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

You want to find its derivative ??

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

See here: \[\large \frac{d}{dx}(\sin(f(x)) = \cos(f(x)) \times \frac{d}{dx}(f(x))\]

OpenStudy (anonymous):

So here: What is the term in place of f(x) in sin(f(x)) ??

OpenStudy (anonymous):

sin has its derivative as cos.. So we are left with the derivative of f(x) That will be: \[\frac{d}{dx}(\sin(\frac{x+1}{2x-3})) = \cos(\frac{x+1}{2x-3}) \times \frac{d}{dx}(\frac{x+1}{2x-3})\]

OpenStudy (anonymous):

so by using division rule you can find the derivative of last term..

OpenStudy (anonymous):

no,,

OpenStudy (anonymous):

\[\large \frac{u}{v} = \frac{v \cdot u' - u \cdot v'}{v^2}\]

OpenStudy (anonymous):

\[\frac{d}{dx}(\frac{x+1}{2x-3}) = \frac{(2x-3)(1) - (x+1)(2)}{(2x-3)^2} \implies \frac{-5}{(2x-3)^2}\]

OpenStudy (anonymous):

Now combine them: \[\frac{d}{dx}(\sin(\frac{x+1}{2x-3})) = \cos(\frac{x+1}{2x-3}) \times \frac{-5}{(2x-3)^2}\]

OpenStudy (australopithecus):

f(x)=sin(x+1/2x−3) use chain rule g(x) = sin(x) g'(x) = cos(x) f(x) = (x+1)/(2x-3) f'(x) = (see below) Deriving f'(x): s(x) = x+1 s'(x) = 1 m(x) = 2x-3 m'(x) = 2 we use \[\frac{s'(x)m(x) - s(x)m'(x)}{(m(x))^2}\] so we have f'(x) = \[\frac{(2x-3) - (2)(x+1)}{(2x-3)^2}\] Now we just use chain rule g(x) = sin(x) g'(x) = cos(x) f(x) = (x+1)/(2x-3) f'(x) = \[\frac{(2x-3) - (2)(x+1)}{(2x-3)^2}\] So knowing this we apply g'(f(x)) * f'(x) so \[\cos(\frac{x+1}{2x-3})*(\frac{(2x-3) - (2)(x+1)}{(2x-3)^2})\]

OpenStudy (australopithecus):

the bottom equation is f'(x), also I made a mistake and used f(x) in the solution I should have used like l(x) or something for (x+1)/(2x-3) so just pretend f(x) is l(x) so, l(x) = (x+1)/(2x-3) to avoid confusion

OpenStudy (australopithecus):

the best way to deal with derivatives when you first start out I find is to just split them up into separate functions, take the derivative of those smaller functions and then stick them back into the applicable formula to get the final answer

OpenStudy (australopithecus):

also note i split l(x) into two functions so l(x) = s(x)/m(x)

OpenStudy (australopithecus):

if you have any questions feel free to ask, really once you do it this way derivatives are super easy eventually you won't even need to split them up, you can just do them in your head super quick

OpenStudy (anonymous):

can u give me all formulas the derived

OpenStudy (australopithecus):

second page of this document

OpenStudy (anonymous):

ok thanks

OpenStudy (australopithecus):

to be honest you don't have to memorize quotient rule, you can just use product rule with chain rule \[\frac{(x+1)}{(2x-1)} = (x+1)(2x-1)^{-1}\]

OpenStudy (australopithecus):

but that is just a note for the future :)

OpenStudy (anonymous):

uv=v⋅u′−u⋅v′v2 u means??

OpenStudy (australopithecus):

you would just have q(x) = x+1 q'(x) = 1 d(x) = (2x - 1)^(-1) use chain rule here d'(x) = -1(2x-1)^(-2) * 2 then just plug into product rule q'(x)d(x) + d'(x)q(x)

OpenStudy (australopithecus):

but if you are just starting with derivatives its probably best you learn quotient

OpenStudy (anonymous):

hey! put your equation in 'WOLFramalpha' and u can get the result its an outstanding computational website just put ur equation in its search engine and it will generate output of ur equation!!!!@Muskan

OpenStudy (anonymous):

http://www.wolframalpha.com/

OpenStudy (australopithecus):

yeah wolfram alpha can be good to check your work but sometimes your answers can end up looking different even though they are correct and the same so just be aware

OpenStudy (anonymous):

@Australopithecus i put above equation in it and it give a correct answer with graph i love it(wolfram alpha)

OpenStudy (australopithecus):

you should advertise that site for a living

OpenStudy (anonymous):

@Australopithecus it makes work load easy!

OpenStudy (anonymous):

it is a nys site thanks @ali110

OpenStudy (anonymous):

fundamental knowledge is the key to solving "difficult" questions